# Investigating the effects of QCD matter’s electrical conductivity on charge dependent directed flow

### v1

@@ VI.1 Centrality dependence of charge dependent flow
In our previous work [21], $\Delta(v_{1})$ of pions were studied instead of protons. The result of that work proposed the electric conductivity of QGP should be $\sim\mathcal{O}(10^{-3})$ or smaller when fit to experimental data. However, experimentally, the values of $\Delta(v_{1})$ depend on the type of hadron (pions vs. protons) and the centrality of the collision [16, 18]. The centrality dependence occurs because the EM field intensity is proportional to the number of collision spectators, which means ultra-central collisions should have a smaller intensity than peripheral collisions. Then the decrease in EM intensity translates to smaller slopes of $\Delta v_{1}$. This trend is reflected in the STAR data of $\Delta(dv_{1}/dy)$ for protons and anti-protons between $20\%$ and $65\%$ centrality [16].
Figure 6 illustrates that our RRMHD model can capture how smaller centralities translate into a smaller $\Delta(dv_{1}/dy)$. The dependence on electrical conductivity changes non-trivially across centralities, with a conductivity similar to lattice $\sigma=0.0294\text{ fm}^{-1}$ having the least change over centralities [7]. In addition, the slope is always negative, which indicates the Faraday current always has a greater net effect than the Lorentz current in our model. However, as was discussed in Ref. [15], the initial protons in the collision mean the plasma would have a net positive charge. Until this point, we have considered the initial charge density to be zero, $q_{0}=0$. The issue is initial dynamics of QGP have not been agreed upon, so a quantitative charge density distribution is difficult to establish. If we consider a phenomenological model like Ref. [37] for the baryon charge density and multiply it by the initial nuclei $n_{Q}/n_{B}$ to calculate the electric charge density, we can get a rough idea of the initial electric charge density $q_{0}$. See appendix A for a detailed discussion on the implementation.
@@
This demonstrates that an initial positive charge can result in positive slopes for $\Delta v_{1}$ with our RRMHD model. This result is dependent on how the initial charge density is defined and the conductivity of QGP. For example, the slope does not become positive for the largest conductivity value in Fig. 6 of $\sigma=0.0294\text{ fm}^{-1}$,
@@
Figure 6: The slope of the difference in charge dependent directed flow for protons-antiprotons across collision centralities. From physical arguments, in central symmetric collisions the magnetic field intensity is weaker, leading to smaller charged currents in the plasma. This is illustrated in our relativistic resistive magneto-hydrodynamic (RRHMD) model’s predictions of a small slopes $\Delta(dv_{1}/dy)$ near 2% centrality and larger slopes as centrality increases. At more peripheral collisions, our model requires a large electric conductivity $\sigma$ to predict similar slope values as STAR [16].
@@ VII Conclusions
In future work, we aim to improve our model in several ways. Our results suggest that a complete modeling of $\Delta(dv_{1}/dy)$ requires a combination of baryon and electric charges. This means incorporating the Transported quark effects of Refs. [37, 38] with the electric conductivity effects of our RRMHD model. It is possible to accomplish that incorporation using a phenomenological approach as described in appendix A, or a more sophisticated model of the initial collision stages. Moreover, a dedicated analysis of how different models of the initial stages change the initial EM fields + charge density would be interesting on its own.
@@
The previous improvements focus on improving the accuracy of calculating the charge density $q$. For a consistent calculation, we should use an equation of state (EoS) that includes the electric chemical potential, for example [39]. Although, considering in this work the charge density was found to be small on the freezeout hypersurface, the impact of changing the EoS to include the electric chemical potential may also be small.
In this work, we have assumed the electric conductivity of QGP is a scalar. But, dimensional analysis shows the conductivity should be temperature-dependent and dedicated calculations of the conductivity agree [7]. Additionally, for intense magnetic fields, the conductivity splits into a component parallel with the magnetic fields and a component perpendicular. Considering both of though effects would change the conductivity from a scalar to a tensor. A future investigation about how a tensor conductivity could affect experimental observables maybe interesting.
We have chosen a single freezeout temperature and then used the Cooper-Frye formula to calculate the final particle distributions. This is quite an oversimplification of the final hadron dynamics that would be better described with an afterburner + EM field calculation. Possible candidates are the kinetic solvers UrQMD, SMASH, or JAM, and ideally, we would like our model to be able to interface with all three.
Lastly, it has been well established that viscous hydrodynamics is required to explain the elliptical flow $v_{2}$ of final state hadrons. Although, it may seem natural to include viscosity in our model, because of how it could modify the Lorentz current, we do not believe it to be as important as other improvements. For example, our results agree with Ref. [15] even though they use iEBE-VISHNU + a perturbative EM field calculation.
@@ Acknowledgements.
We thank Azumi Sakai for useful discussions and a thorough review of the initial manuscript. Furthermore, we thank Hidetoshi Taya and Akihiko Monnai for insightful discussions about early versions of this work. The numerical computation in this work was carried out at the Yukawa Institute Computer Facility. N.J.B. is grateful for the financial support from Hiroshima University’s presidential discretionary fund and Start-Up program fund. This work was also supported by JSPS KAKENHI Grant Numbers, JP20K11851, JP20H00156, JP24K07117 (T.M.), JP20H00156, JP20H11581 (C.N.), JP21H04488, JP24K00672, and JP24K00678 (H.R.T.). Perceptually uniform color maps are used in this study to prevent visual distortion of the data and to better serve those with color vision deficiency [40].
Appendix A Effects from an initial charge density
@@
where $T(x\pm b/2,y)$ are the normalized nuclei Fermi distributions. We have used a radius of $r_{0}=6.38$, a width of $\delta=0.535$, and a nuclear saturation density of $\rho_{0}=0.17$ for 197Au. The functions $f_{\pm}(\eta)$ in Eq.(23) impose a tilt in the longitudinal direction and $H(\eta)$ define a longitudinal profile [31]. For the initial baryon number calculation in Ref. [37], an additional longitudinal weighting is preformed with $H_{B}(\eta)$ that has two free parameters $\eta_{B}$ and $\sigma_{B}$. We ignore $H_{B}(\eta)$ in this work because fixing those parameters requires solving the hydrodynamic evolution with a nonzero baryon current such that the ratio of transported and produced baryons at central rapidities can be reproduced. We do include a ratio parameter $Z_{q}$ between the initial number of baryons vs. number of protons, which acts as a conversion from the initial baryon charge to the initial electric charge.
Appendix B Charge dependent freeze-out algorithm
We start from the usual Cooper-Frye formula,
[…]
and take the local thermal distribution,
where $Q$ is the quantized electric charge of the hadron. For this work, we simplify the electric charge chemical potential $\mu_{Q}(x)$ to the linear approximation of the local charge density $q(x)$. This results in,
$\mu_{Q}(x)\simeq\frac{q(x)}{g_{h}T_{f}^{2}}$ (27)
where $g_{h}$ is the hadronic degrees of freedom.
Appendix C Dependence on the intensity of the initial electromagnetic fields
We mentioned in Sec. III the pre-equilibrium dynamics after the collision could change the configuration of the initial EM fields. Results in Ref. [26], which uses a Kinetic model for the pre-equilibrium dynamics, suggests the EM fields decay like in a vacuum initially. The initial condition we used assumes the EM fields are in a conducting medium, thus the EM fields decay is that of in a medium [25]. After comparing the two methods at the initialization time of our simulation $t_{0}=0.4\text{ fm}/c$, we find a difference of approximately $0.1$ in the intensity of the magnetic field $B_{y}$ at the center of the grid.
We reproduce the $\sigma=0.0294\text{ fm}^{-1}$ panel of Fig. 5 after decreasing the intensity of initial EM fields by 0.1 across the entire computational domain. In Fig. 7, the slope of $\Delta v_{1}$ is reduced by an order of magnitude proportional to the same reduction of the initial EM fields from $\mathcal{O}(10^{-4})$ to $\mathcal{O}(10^{-5})$. This suggests our results are equally sensitive to the electric conductivity of QGP $\sigma$ and the initial EM fields.
Figure 7: Charge dependent directed flow, calculated using our relativistic resistive magneto-hydrodynamic (RRMHD) model [22]. Compared to panel b of Fig.5, the initial EM fields have been reduced by a multiplication factor of $0.1$. The resulting $\Delta v_{1}$ slope also decreases by a factor $0.1$.

### v2

@@ VI.1 Centrality dependence of charge dependent flow
In our previous work [21], $\Delta(v_{1})$ of pions were studied instead of protons. The result of that work proposed the electric conductivity of QGP should be $\sim\mathcal{O}(10^{-3})$ or smaller when fit to experimental data. However, experimentally, the values of $\Delta(v_{1})$ depend on the type of hadron (pions vs. protons) and the centrality of the collision [16, 18]. The centrality dependence occurs because the EM field strength is proportional to the number of collision spectators, which means ultra-central collisions should have a smaller strength than peripheral collisions. Then the decrease in the EM strength translates to smaller slopes of $\Delta v_{1}$. This trend is reflected in the STAR data of $\Delta(dv_{1}/dy)$ for protons and anti-protons between $20\%$ and $65\%$ centrality [16].
Figure 6 illustrates that our RRMHD model can capture how smaller centralities translate into a smaller $\Delta(dv_{1}/dy)$. The dependence on electrical conductivity changes non-trivially across centralities, with a conductivity similar to lattice $\sigma=0.0294\text{ fm}^{-1}$ having the least change over centralities [7]. In addition, the slope is always negative, which indicates the Faraday current always has a greater net effect than the Lorentz current in our model. However, as was discussed in Ref. [15], the initial protons in the collision mean the plasma would have a net positive charge. Until this point, we have considered the initial charge density to be zero, $q_{0}=0$. The issue is initial dynamics of QGP have not been agreed upon, so a quantitative charge density distribution is difficult to establish. If we consider a phenomenological model like Ref. [40] for the baryon charge density and multiply it by the initial nuclei $n_{Q}/n_{B}$ to calculate the electric charge density, we can get a rough idea of the initial electric charge density $q_{0}$. See Appendix B for a detailed discussion on the implementation.
@@
This demonstrates that an initial positive charge can result in positive slopes for $\Delta v_{1}$ with our RRMHD model. The possibility of a positive slope also depends on the conductivity of QGP. For example, the slope does not become positive for the largest conductivity value in Fig. 6 of $\sigma=2.94\text{ fm}^{-1}$,
@@
Even though we started with an initial electric charge density $q_{0}\neq 0$.
Figure 6: The slope of the difference in charge dependent directed flow for protons-antiprotons across collision centralities. From physical arguments, in central symmetric collisions the magnetic field strength is weaker, leading to smaller charged currents in the plasma. This is illustrated in our relativistic resistive magneto-hydrodynamic (RRHMD) model’s predictions of a small slopes $\Delta(dv_{1}/dy)$ near 2% centrality and larger slopes as centrality increases. At more peripheral collisions, our model requires a large electric conductivity $\sigma$ to predict similar slope values as STAR [16].
@@ VII Conclusions
In future work, we aim to improve our model in several ways. Our results suggest that a complete modeling of $\Delta(dv_{1}/dy)$ requires a combination of baryon and electric charges. This means incorporating the transported quark effects of Refs. [40, 41, 42, 43] with the electric conductivity effects of our RRMHD model. It is possible to accomplish that incorporation using a phenomenological approach as described in Appendix B, or a more sophisticated model of the initial collision stages. Moreover, a dedicated analysis of how different models of the initial stages change the initial EM fields + charge density would be interesting on its own.
@@
The previous improvements focus on improving the accuracy of calculating the charge density $q$. For a consistent calculation, we should use an equation of state (EoS) that includes the electric chemical potential during the time-evolution, for example [37]. Considering this work found the charge density to be small on the freezeout hypersurface, see Figs. 3 and 4, the impact of changing the EoS during the time-evolution to include the electric chemical potential may also be small.
In this work, we have assumed the electric conductivity of QGP is a scalar. But, dimensional analysis shows the conductivity should be temperature-dependent and dedicated calculations of the conductivity agree [7]. Additionally, for intense magnetic fields, the conductivity splits into a component parallel with the magnetic fields and a component perpendicular. Considering both of those effects would change the conductivity from a scalar to a tensor. A future investigation about how a tensor conductivity could affect experimental observables maybe interesting.
We have chosen a single freezeout temperature and then used the Cooper-Frye formula to calculate the final particle distributions. This is quite an oversimplification of the final hadron dynamics that would be better described with an afterburner + EM field calculation. Possible candidates are the kinetic solvers UrQMD [44], SMASH [45], or JAM [46], and ideally, we would like our model to be able to interface with all three.
Lastly, it has been well established that viscous hydrodynamics is required to explain the elliptical flow $v_{2}$ of final state hadrons. Although, it may seem natural to include viscosity in our model, because of how it could modify the Lorentz current, we do not believe it to be as important as other improvements. For example, our results agree with Ref. [15] even though they use iEBE-VISHNU [39] + a perturbative EM field calculation.
@@ Acknowledgements.
We thank Azumi Sakai for useful discussions and a thorough review of the initial manuscript. Furthermore, we thank Hidetoshi Taya and Akihiko Monnai for insightful discussions about early versions of this work. The numerical computation in this work was carried out at the Yukawa Institute Computer Facility. N.J.B. is grateful for the financial support from Hiroshima University’s presidential discretionary fund and Start-Up program fund. This work was also supported by JSPS KAKENHI Grant Numbers, JP20K11851, JP20H00156, JP24K07117 (T.M.), JP20H00156, JP20H11581 (C.N.), JP21H04488, JP24K00672, and JP24K00678 (H.R.T.) and by the World Premier International Research Center Initiative (WPI) under MEXT, Japan (C.N.). Perceptually uniform color maps are used in this study to prevent visual distortion of the data and to better serve those with color vision deficiency [47].
Appendix A Initial electromagnetic fields
For the work, we calculated the initial EM fields using the results of Ref. [25] which we briefly summarize. Starting from the EM fields produced by a single charge $e$ only moving along the beam axis $z$ with a fixed velocity $v$, the in-medium Maxwell equations are solved. By taking the relativistic limit $\gamma=1/\sqrt{1-v^{2}}\gg 1$ and accounting for a conductive medium $\gamma\sigma b\gg 1$, the solutions to Maxwell equations become,
$E_{r}=B_{\phi}=\frac{e}{2\pi}\frac{b\sigma}{4(t\pm v/z)^{2}}\exp{\left(-\frac{b^{2}\sigma}{4(t\pm v/z)}\right)},$ (27)
$E_{z}=-\frac{e}{4\pi}\frac{(t\pm v/z)-b^{2}\sigma}{\gamma^{2}(t\pm v/z)^{3}}\exp{\left(-\frac{b^{2}\sigma}{4(t\pm v/z)}\right)},$ (28)
in cylindrical coordinates. Those are the same equations previously used in Ref. [23], but we have suppressed the $\hbar c$ factors and the dimensions of the EM fields should be understood implicitly to be $\sqrt{\text{GeV}/\text{fm}^{3}}$.
We account for the multiple charges of heavy-ions like 197Au in Tab. 1 by diffusing the total charge equally inside the ion $\rho=Z/\left(\frac{4}{3}\pi R^{3}\right)$. Here $Z$ is the number of protons inside the ion and $R$ is the ion’s radius. Then we integrate over the contribution of all the charge densities to arrive at the EM field configuration created by a single ion moving along the $z$ axis, which is step 2 of Sec. IV. The same method is applied to the other ion involved in the collision, but the sign of the velocity is reversed v-v [25].
Because the charge is equally distributed inside the ion, we refer to this method as smooth initial conditions. This is contrary to fluctuating initial conditions, where the charge density can vary inside the ion.
Appendix B Effects from an initial charge density
@@
where $T(x\pm b/2,y)$ are the normalized nuclei Fermi distributions and $A$ is the nuclear mass. For this work, we have used a radius of $r_{0}=6.38\text{ fm}$, a width of $\delta=0.535\text{ fm}$, and a nuclear saturation density of $\rho_{0}=0.17\text{ fm}^{-3}$ for 197Au. The functions $f_{\pm}(\eta)$ in Eq.(29) impose a tilt in the longitudinal direction and $H(\eta)$ define a longitudinal profile [31]. For the initial baryon number calculation in Ref. [40], an additional longitudinal weighting is preformed with $H_{B}(\eta)$ that has two free parameters $\eta_{B}$ and $\sigma_{B}$. We ignore $H_{B}(\eta)$ in this work because fixing those parameters requires solving the hydrodynamic evolution with a nonzero baryon current such that the ratio of transported and produced baryons at central rapidities can be reproduced. We do include a ratio parameter $Z_{q}$ between the initial number of baryons vs. number of protons, which acts as a conversion from the initial baryon charge to the initial electric charge.
Appendix C Dependence on the strength of the initial electromagnetic fields
We mentioned in the second bullet point of Sec. III the pre-equilibrium dynamics after the collision could change the configuration of the initial EM fields. An example of a change is the results in Ref. [26], which uses a Kinetic model for the pre-equilibrium dynamics. Their results suggest the EM fields initially decay with a time dependence of $\lvert eB\rvert\sim\tau^{-3}$, like in a vacuum [48]. Then the decay smoothly transition to a time dependence of $\lvert eB\rvert\sim\tau^{-1}$, which is an in-medium like decay behavior [3, 49]. In contrast, the initial condition we used assumes the EM fields always decay in a conducting medium [25]. This indicates our initial conditions overestimate the strength of the EM fields.
Figure 7: Charge dependent directed flow, calculated using our relativistic resistive magneto-hydrodynamic (RRMHD) model [22]. Compared to panel b of Fig.5, the initial EM fields have been reduced to $10\%$ of the original values. All other parameters are the same as panel b. The resulting $\Delta v_{1}$ slope also decreased on the order of $10\%$ relative to the slope in panel b.
To roughly understand how that overestimation could affect our results, we have scaled down the strength of the initial EM fields to $10\%$ of the original values [25]. Then we reproduced panel b) of Fig. 5, for which the conductivity was similar to lattice QCD values $\sigma=0.0294\text{ fm}^{-1}$. The result is illustrated in Fig. 7, the slope of $\Delta v_{1}$ becomes reduced by an order of magnitude from $\mathcal{O}(10^{-4})$ to $\mathcal{O}(10^{-5})$. This is approximately proportional to the same reduction of the initial EM fields of $10\%$. Although this suggests our results are equally sensitive to the electric conductivity of QGP $\sigma$ and the initial EM field intensities, more robust modeling of the pre-equilibrium dynamics is required to draw any firm conclusions.